The function, p(d)=1+d/33, gives the pressure, in atmospheres (atm), at a depth d in the sea (d is in feet). Note that p(0)=1 atm, p(33)=2, and so on. Find the pressure at 150 ft.
The pressure at 150 feet is ____ atm.

Answer by nerdybill(7384)

The problem gives you a function of depth:
p(d)=1+d/33
where d = depth in feet
.
And, if d = 150 ft
then
p(150)=1+150/33
p(150)=1+4.545
p(150)=5.545 atm
.
Therefore:
The pressure at 150 feet is 5.545 atm.